Suppose you are using a=0.05 to test the claim that u=25 using a p-value. sample statistics n=35, x=24.1, and o=2.7 Find the p-value
Answer by stanbon(75887)

Suppose you are using a=0.05 to test the claim that u=25 using a p-value. sample statistics n=35, x=24.1, and o=2.7 Find the p-value
-------------
2-tail test
Using a T-test:
t(24.1) = (24.1-25)/[2.7/sqrt(35)] = -1.9720
---
p-value = 2*P(-100< t < -1.9720, when df = 34)
= tcdf(-100,-1.9720,34)
= 0.028395
=============
Cheers,
Stan H.